40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jan 23 23:10:19 2014, ericq , Update, General, Vent so far 
    Reply  Mon Jan 27 21:36:35 2014, ericq , Update, General, Vent so far 
       Reply  Tue Jan 28 22:24:46 2014, ericq , Update, General, Further Alignment 
Message ID: 9579     Entry time: Mon Jan 27 21:36:35 2014     In reply to: 9572     Reply to this: 9583
Author: ericq  
Type: Update 
Category: General 
Subject: Vent so far 

After turning the slow FSS threshold down, the mode cleaner stays locked enough to do other things. We were able to align the tip tilts to the y-arm such that we were able to get some flashes in what looks like a TM00-ish mode. (It was necessary to align the PRM such that there was some extra power circulating in the PRC to be able to see the IR flashes on the ITMY face camera) This is enough to convince us that we are at least near a reasonable alignment, even though we couldn't lock to the mode. 

The x-arm was in a hairier situation; since the green beam wouldn't flash into any modes, we don't even know that a good cavity axis exists. So, I used the green input PZTs to shine the green beam directly on the earthquake stops on the ITMX cage, and then inferred the PZT coordinates that would place the green beam roughly on the center of ITMX. I moved the ETMX face camera such that it points at the ETMX baffle. I tried looking for the retroreflected green spot to no avail. Hopefully tomorrow, we can get ourselves to a reasonably aligned state, so we can begin measuring the macroscopic PRC length. 

ELOG V3.1.3-